Find the sumof 2 consecutive positive even integers such that the square of the smaller is 4 more than 10 times the larger.
Help!

Answer by mananth(16946)

let the integers be x & x+2
...
x2=10(x+2)+4
x^2=10x+20+4
x^2=10x+24
x^2-10x-24=0
x^2-12x+2x-24=0
x(x-12)+2(x-12)=0
(x-12)(x+2)=0
x=12 OR x=-2( ignore)
the integers are 12 14
...
m.ananth@hotmail.ca